# SHREC: A Framework for Advancing Next-Generation Computational Phenotyping with Large Language Models

### v2

@@
Abstract
Objective: Computational phenotyping is a central informatics activity with resulting cohorts supporting a wide variety of applications. However, it is time-intensive because of manual data review, limited automation, and difficulties in adapting algorithms across sources. Since LLMs have demonstrated promising capabilities for text classification, comprehension, and generation, we posit they will perform well at repetitive manual review tasks traditionally performed by human experts. To support next-generation computational phenotyping methods, we developed SHREC, a framework for comprehensive integration of LLMs into end-to-end phenotyping pipelines.
Methods: We applied and tested three lightweight LLMs (Gemma billion, Mistral Small billion, and Phi- billion) to classify concepts and phenotype patients using previously developed phenotypes for ARF respiratory support therapies.
Results: All models performed well on concept classification, with the best model (Mistral) achieving an AUC of across all relevant concepts. For phenotyping, models demonstrated near-perfect specificity for all phenotypes, and the top-performing model (Mistral) reached an average AUC of for single-therapy phenotypes, despite lower performance on multi-therapy phenotypes.
Conclusion: Current lightweight LLMs can feasibly assist researchers with resource-intensive phenotyping tasks such as manual data review. There are several advantages of LLMs that support their application to computational phenotyping, such as their ability to adapt to new tasks with prompt engineering alone and their ability to incorporate raw EHR data. Future steps to advance next-generation phenotyping methods include determining optimal strategies for integrating biomedical data, exploring how LLMs reason, and advancing generative model methods.
keywords:
Large Language Models , Generative Artificial Intelligence , Computational Phenotyping , Electronic Phenotyping[label1] organization=Program in Applied Mathematics, Department of Mathematics, The University of Arizona, city=Tucson, state=AZ, postcode=85721, country=United States
[label2] organization=College of Engineering, The University of Arizona, city=Tucson, state=AZ, postcode=85721, country=United States
[label3] organization=College of Medicine - Tucson, The University of Arizona, city=Tucson, state=AZ, postcode=85721, country=United States
1 Introduction
@@
| Problem or Issue | What is Already Known | What this Paper Adds | Who Would Benefit |
| Computational phenotyping is fundamental, yet time and resource-intensive due to manual data review and limited scalability. | Traditional approaches, including advanced Natural Language Processing and machine learning methods, require significant development and lack generalizability across clinical settings, indicating the need for novel, next-generation phenotyping methods. | This study introduces SHREC (SHifting to language model-based REal-world Computational phenotyping), a framework for integrating LLMs into phenotyping pipelines. We demonstrated lightweight LLMs can accurately classify concepts and phenotype encounters using raw EHR data and prompt engineering alone. This is the first study to develop LLM-based methods for this application. | Biomedical informaticians and clinical researchers seeking more efficient approaches to computational phenotyping and cohort identification. |
2 Related Work
3 Methods
@@
3.1 Development of SHREC
3.1.1 Theoretical Foundation
@@ 3.1.2 Framework Components
3.2 Application of SHREC to Phenotyping Use Case
3.2.1 Phenotyping Use Case
@@
3.2.2 Identification of Phenotyping Tasks
In the original study, data from the eICU Collaborative Research Database (eICU-CRD)[26] were manually reviewed to first determine relevance to the therapies or medications of interest and second to produce a phenotyping algorithm.[27] These processes mapped to the Concept Selection and Apply Algorithm tasks of the end-to-end phenotyping pipeline within SHREC, respectively. In this study, we used LLMs to perform the tasks of Concept Selection and Apply Algorithm while we also implemented Outline Phenotype Heuristics, LLM Model Selection, and Evaluate Algorithm because they were required to execute and test the LLM-based methods. Since the remaining tasks were not related to the LLM methods, they were not included in this study, but were previously discussed in-depth.[27, 12]
3.3 Implementation of LLMs for Phenotyping Tasks
@@
3.3.1 Outline Phenotype Heuristics
@@
3.3.2 LLM Model Selection
@@ 3.3.3 Concept Selection
For Concept Selection, we generated constructed concepts from eICU-CRD tables. Tables that did not include timestamped data but contained information on respiratory therapies or medications (such as the apacheApsVar table) or were unlikely to contain descriptions of respiratory therapies or medications (such as the vitalPeriodic table) were not processed further. There were tables used to construct input concepts (Table 3). In early testing, we achieved best results when classifying the respiratory therapies separately from the relevant medications. We thus developed two prompts, resulting in two LLM responses per constructed concept (see Supplementary Material). Concept definitions within the prompts were produced by two clinician experts who summarized and consolidated notes related to the terms of interest (see Supplementary Material). We used Chain-of-Thought (CoT)prompting by including a series of questions and answers to help the model determine relevancy of the constructed concept. CoT is a prompt engineering technique that has improved performance of LLMs by using a series of reasoning tasks to guide the model to the final response.[30, 31] The answer to the final question for each prompt was parsed using string methods to get the final response of “YES” or “NO” for whether the constructed concept was relevant. Since there were two prompts, the final response was “YES” if at least one of the responses was “YES” and “NO” otherwise.
3.3.4 Apply Algorithm
@@
After we identified the relevant constructed concepts across the eICU-CRD dataset, we applied the phenotyping heuristics. We identified the first encounter for each unique patient and then removed individuals who were less than years old at the start of the encounter. For each encounter, we filtered data from the original tables to the selected constructed concepts and then ordered each distinct constructed concept by its first occurrence based on the encounter admission time. The data, or constructed descriptions, for phenotyping were created by inserting each individual constructed concept into a string template and concatenating all of the unique constructed concepts together for the prompt. The template was “#: {constructed concept}” where “#” was the order of the constructed concept based on its first occurrence in the encounter records. Since timestamps were generalized to concept order and there was no encounter-specific information in the constructed descriptions, we phenotyped the unique constructed descriptions and then mapped the phenotypes to the relevant encounters for evaluation. Our phenotyping prompt used CoT (see Supplementary Material). We parsed the answer to the final question to identify the selected phenotype.
3.3.5 Evaluate Algorithm
@@
Models were evaluated for both Concept Selection and Apply Algorithm using components of PHEONA, an evaluation framework for LLM-based applications to computational phenotyping.[18] Previously, we evaluated the models for Concept Selection using a random sample of constructed concepts.[18] In this study, we evaluated these models on the full set of constructed concepts using Accuracy (the ability of the model to produce accurate results) as the primary evaluation criterion and Model Response Latency (how quickly model results were returned) as the secondary evaluation criterion from PHEONA. We used area under the receiver operating characteristic curve (AUC) to measure response accuracy against the concept ground truths. We measured response latency as the seconds required for the model to return a response for each constructed concept and then averaged these values for each prompt. For Apply Algorithm, since we had not previously used PHEONA for model evaluation, we evaluated model performance on a randomly selected subsample (see Supplementary Material) and then evaluated the best performing models on all encounters. We also used Accuracy and Model Response Latency criteria to assess model performance with the AUC (and additionally, sensitivity and specificity) calculated using the original encounter ground truths.[27]
4 Results
4.1 Phenotyping Use Case
@@
4.2 Implementation of LLMs for Phenotyping Tasks
4.2.1 Concept Selection
There were concept ground truths ( ARF respiratory support therapies and medications) from the original phenotyping study.[27] Classification results based on the concept ground truths are presented in Table 4. Mistral had the highest accuracy with an AUC of for classification of all concepts; however, it also had the highest total average latency ( seconds compared to seconds and faster). All models performed better at medication classification (AUC of and higher) when compared to respiratory support therapy classification ( and higher).
4.2.2 Apply Algorithm
There were unique constructed descriptions for Gemma, for Mistral, and for Phi. However, since Gemma underperformed on the subsample of constructed descriptions (see Supplementary Material), only Mistral and Phi were tested on the entire dataset. Mistral had an average response latency of seconds and Phi of seconds across all constructed descriptions. The AUC, sensitivity, and specificity for each phenotype are presented in Table 5. Overall, Mistral performed better with a higher AUC on all phenotypes when compared to Phi. Mistral performed best on no therapy and single therapy phenotypes (None and IMV, NIPPV, and HFNI Only) with an average AUC of while it only achieved an average AUC of on the remaining, multi-therapy phenotypes. Both models also had nearly perfect specificity for all phenotypes except IMV Only.
5 Discussion
@@
5.1 Development and Application of SHREC
@@
5.2 Implementation of LLMs for Phenotyping Tasks
@@ 5.3 Study Limitations
5.4 Future Directions
There are many future directions to explore outside of those highlighted by the broader vision of computational phenotyping from SHREC. One direction is to understand how LLMs reason with respect to computational phenotyping. Although we used CoT because of its demonstrated ability to produce accurate results,[30, 31, 32] recent studies have suggested that CoT reasoning may actually be unfaithful.[34, 35, 36] Given the complexity of phenotyping tasks, we suggest studying CoT reasoning to understand when and how logical inconsistencies may arise. Another future direction is to adapt phenotype definitions to LLMs. For example, the clinician experts indicated some of the medications may be present for noninvasive therapies, but not as a continuous infusion (see Supplementary Material). However, for the previously developed definition, only medication presence was used due to complexities in algorithmically processing medication information in relation to respiratory therapies. In future iterations, we would update the phenotype definition to include administration method rather than simply asking the model to look for concept presence. Finally, we propose development of industry standards for evaluation of LLM-based methods specifically with respect to automated processes to ensure appropriate oversight.
6 Conclusion
@@
7 Declaration of Competing Interests
The authors have no competing interests to declare.
8 Funding
This work was supported by the National Institute of General Medical Sciences of the National Institutes of Health under grant T GM.
9 Data Statement
All data used for this study are available through the eICU Collaborative Research Database (eICU-CRD)database (https://eicu-crd.mit.edu/). All code was run with Python version and can be found at our repository (https://github.com/spungit/SHRECandPHEONA).

### v4

@@
SHREC: A Framework for Advancing Next-Generation Computational Phenotyping with Large Language Models
Sarah A. Pungitore, MS1, Shashank Yadav, MS2, Molly Douglas, MD3, Jarrod Mosier, MD3, Vignesh Subbian, PhD2
1 Program in Applied Mathematics, The University of Arizona, Tucson, AZ
2 College of Engineering, The University of Arizona, Tucson, AZ
3 College of Medicine - Tucson, Tucson, AZ
1.1. Abstract
Computational phenotyping is a central informatics activity with resulting cohorts supporting a wide variety of applications. However, it is time-intensive because of manual data review and limited automation. Since LLMs have demonstrated promising capabilities for text classification, comprehension, and generation, we posit they will perform well at repetitive manual review tasks traditionally performed by human experts. To support next-generation computational phenotyping, we developed SHREC, a framework for integrating LLMs into end-to-end phenotyping pipelines. We applied and tested three lightweight LLMs (Gemma billion, Mistral Small billion, and Phi- billion) to classify concepts and phenotype patients using phenotypes for ARF respiratory support therapies. All models performed well on concept classification, with the best (Mistral) achieving an AUROC of . For phenotyping, models demonstrated near-perfect specificity for all phenotypes with the top-performing model (Mistral) achieving an average AUROC of for single-therapy phenotypes. In conclusion, lightweight LLMs can assist researchers with resource-intensive phenotyping tasks. Several advantages of LLMs included their ability to adapt to new tasks with prompt engineering alone and their ability to incorporate raw EHR data. Future steps include determining optimal strategies for integrating biomedical data and understanding reasoning errors.
1.2. Introduction
@@
1.3. Materials and Methods
@@
1.3.1. Development of SHREC
1.3.1.1. Theoretical Foundation
@@ 1.3.1.2. Framework Components
1.3.2. Application of SHREC to Phenotyping Use Case
1.3.2.1. Phenotyping Use Case
@@
Step | Name | Description | Analog to Previous Framework |
1 | Assess Fitness-for-Purpose | Determine the clinical outcome of interest, assess clinical significance, and assess any sources of clinical or data complexity. | Assess Fitness-for-Purpose |
2 | Identify Data Sources | Identify data source(s) to use for phenotype development and evaluation. | Assess Fitness-for-Purpose |
3 | Gather Ground Truth Data | Determine ground truth labels to use for validation of the phenotyping algorithm. | Create Gold Standard Data |
4 | Outline Phenotyping Heuristics | Determine the tasks necessary in the phenotyping process to obtain the resulting phenotypes from the input data. Will likely include the inclusion and exclusion criteria. | None |
5 | LLM Model Selection | If used, determine which LLMs can be tested for specific tasks and how these models can be evaluated.[60] For studies not using LLMs, can either skip or identify other machine learning or Natural Language Processing (NLP)models. | Develop Models |
6 | Concept Selection | Classify or identify relevant data elements from the electronic health record (EHR)data. | Engineer Features |
7 | Apply Algorithm | Apply the algorithm tasks to each record and identify the appropriate phenotype. | None |
8 | Evaluate Algorithm | Determine the effectiveness of the phenotyping algorithm against the ground truth data. | Evaluate Models |
1.3.2.2. Identification of Phenotyping Tasks
A comparison of the methods performed in the original study and this study are presented in Figure 2. In the original study, data from the eICU Collaborative Research Database (eICU-CRD)[59] were manually reviewed to first determine relevance to the therapies or medications of interest and second to produce a phenotyping algorithm.[47] These processes mapped to the Concept Selection and Apply Algorithm tasks of the end-to-end phenotyping pipeline within SHREC, respectively. In this study, we used LLMs to perform the tasks of Concept Selection and Apply Algorithm while we also implemented Outline Phenotype Heuristics, LLM Model Selection, and Evaluate Algorithm because they were required to execute and test the LLM-based methods. Since the remaining tasks were not related to the LLM methods, they were not included in this study, but were previously discussed in-depth.[47, 43]
1.3.3. Implementation of LLMs for Phenotyping Tasks
@@
1.3.3.1. Outline Phenotype Heuristics
@@
Table Name | Constructed Concept Pattern | Example Constructed Concept |
Care Plan General | Source = Care Plan General; Concept = cplgroup: cplitemvalue | Source = Care Plan General; Concept = Route-Status: Oral - low sodium |
Infusion Drug | Source = Infusion Drug; Concept = drugname | Source = Infusion Drug; Concept = Amiodarone (mg/min) |
Medication | Source = Medication; Concept = drugname | Source = Medication; Concept = LOPRESSOR |
Note | Source = Note; Concept = notevalue: notetext | Source = Note; Concept = denies fevers: denies fevers |
Nurse Care | Source = Nurse Care; Concept = cellattributevalue | Source = Nurse Care; Concept = emergency equipment at bedside |
Nurse Charting | Source = Nurse Charting; Concept = nursingchartcelltypevalname: nursingchartvalue | Source = Nurse Charting; Concept = O2 Admin Device: BiPAP/CPAP |
Respiratory Care | Source = Respiratory Care; Concept = airwaytype | Source = Respiratory Care; Concept = Oral ETT |
Respiratory Charting | Source = Respiratory Charting; Concept = respcharttypecat: respchartvaluelabel: respchartvalue | Source = Respiratory Charting; Concept = respFlowPtVentData: SaO2: 25 |
Treatment | Source = Treatment; Concept = treatmentstring | Source = Treatment; Concept = cardiovascular—myocardial ischemia / infarction—antiplatelet agent—aspirin |
1.3.3.2. LLM Model Selection
@@ 1.3.3.3. Concept Selection
For Concept Selection, we generated constructed concepts from eICU-CRD tables. Tables that did not include timestamped data but contained information on respiratory therapies or medications (such as the apacheApsVar table) or were unlikely to contain descriptions of respiratory therapies or medications (such as the vitalPeriodic table) were not processed further. There were tables used to construct input concepts (Table 2). In early testing, we achieved best results when classifying the respiratory therapies separately from the relevant medications. We thus developed two prompts, resulting in two LLM responses per constructed concept (Supplementary Material). Concept definitions within the prompts were produced by two clinician experts who summarized and consolidated notes related to the terms of interest (Supplementary Material). We used Chain-of-Thought (CoT)prompting by including a series of questions and answers to help the model determine relevancy of the constructed concept. CoT is a prompt engineering technique that has improved performance of LLMs by using a series of reasoning tasks to guide the model to the final response.[64, 68] The answer to the final question for each prompt was parsed using string methods to get the final response of “YES” or “NO” for whether the constructed concept was relevant. Since there were two prompts, the final response was “YES” if at least one of the responses was “YES” and “NO” otherwise.
1.3.3.4. Apply Algorithm
@@
After we identified the relevant constructed concepts across the eICU-CRD dataset, we applied the phenotyping heuristics. We identified the first encounter for each unique patient and then removed individuals who were less than years old at the start of the encounter. For each encounter, we filtered data from the original tables to the selected constructed concepts and then ordered each distinct constructed concept by its first occurrence based on the encounter admission time. The data, or constructed descriptions, for phenotyping were created by inserting each individual constructed concept into a string template and concatenating all of the unique constructed concepts together for the prompt. The template was “#: {constructed concept}” where “#” was the order of the constructed concept based on its first occurrence in the encounter records. Since timestamps were generalized to concept order and there was no encounter-specific information in the constructed descriptions, we phenotyped the unique constructed descriptions and then mapped the phenotypes to the relevant encounters for evaluation. Our phenotyping prompt used CoT (Supplementary Material). We parsed the answer to the final question to identify the selected phenotype.
1.3.3.5. Evaluate Algorithm
@@
Models were evaluated for both Concept Selection and Apply Algorithm using components of PHEONA, an evaluation framework for LLM-based applications to computational phenotyping.[60] Previously, we evaluated the models for Concept Selection using a random sample of constructed concepts.[60] In this study, we evaluated these models on the full set of constructed concepts using Accuracy (the ability of the model to produce accurate results) as the primary evaluation criterion and Model Response Latency (how quickly model results were returned) as the secondary evaluation criterion from PHEONA. We used area under the receiver operating characteristic curve (AUROC) to measure response accuracy against the concept ground truths. We measured response latency as the seconds required for the model to return a response for each constructed concept and then averaged these values for each prompt. For Apply Algorithm, since we had not previously used PHEONA for model evaluation, we evaluated model performance on a randomly selected subsample (Supplementary Material) and then evaluated the best performing models on all encounters. We also used Accuracy and Model Response Latency criteria to assess model performance with the AUROC (and additionally, sensitivity and specificity) calculated using the original encounter ground truths.[47]
1.4. Results
1.4.1. Phenotyping Use Case
@@
1.4.2. Implementation of LLMs for Phenotyping Tasks
1.4.2.1. Concept Selection
There were concept ground truths ( ARF respiratory support therapies and medications) from the original phenotyping study.[47] Classification results based on the concept ground truths are presented in Table 3. Mistral had the highest accuracy with an AUROC of for classification of all concepts; however, it also had the highest total average latency ( seconds compared to seconds and faster). All models performed better at medication classification (AUROC of and higher) when compared to respiratory support therapy classification ( and higher).
Model | Total Concepts |
N | AUROCa | Latency | N | AUROCa | Latency | N | AUROCa | Latency | |||||
Gemma | |||||||||||||
Mistral | |||||||||||||
Phi | |||||||||||||
a AUROC: Area under the receiver operating characteristic curve. |
1.4.2.2. Apply Algorithm
There were unique constructed descriptions for Gemma, for Mistral, and for Phi. However, since Gemma underperformed on the subsample of constructed descriptions (Supplementary Material), only Mistral and Phi were tested on the entire dataset. Mistral had an average response latency of seconds and Phi of seconds across all constructed descriptions. The AUROC, sensitivity, and specificity for each phenotype are presented in Table 4. Overall, Mistral performed better with a higher AUROC on all phenotypes when compared to Phi. Mistral performed best on no therapy and single therapy phenotypes (None and IMV, NIPPV, and HFNI Only) with an average AUROC of while it only achieved an average AUROC of on the remaining, multi-therapy phenotypes. Both models also had nearly perfect specificity for all phenotypes except IMV Only.
1.5. Discussion
@@
1.5.1. Development and Application of SHREC
@@
1.5.2. Implementation of LLMs for Phenotyping Tasks
@@
Phenotype | N | Mistral | Phi | ||||||
N | AUROCa | Sens.b | Spec.c | N | AUROCa | Sens.b | Spec.c | ||
IMV d Only | |||||||||
NIPPV e Only | |||||||||
HFNI f Only | |||||||||
NIPPV e Failure | |||||||||
HFNI f Failure | |||||||||
IMV d to NIPPV e | |||||||||
IMV d to HFNI f | |||||||||
None | |||||||||
a AUROC: Area under the receiver operating characteristic curve. | |||||||||
b Sens.: Sensitivity | |||||||||
c Spec.: Specificity | |||||||||
d IMV: Invasive Mechanical Ventilation. | |||||||||
e NIPPV: Noninvasive Positive Pressure Ventilation. | |||||||||
f HFNI: High-Flow Nasal Insufflation. |
@@ 1.5.3. Study Limitations
1.5.4. Future Directions
There are many future directions to explore outside of those highlighted by the broader vision of computational phenotyping from SHREC. One direction is to understand how LLMs reason with respect to computational phenotyping. Although we used CoT because of its demonstrated ability to produce accurate results,[64, 68, 72] recent studies have suggested that CoT reasoning may actually be unfaithful.[45, 38, 67] Given the complexity of phenotyping tasks, we suggest studying CoT reasoning to understand when and how logical inconsistencies may arise. Another future direction is to adapt phenotype definitions to LLMs. For example, the clinician experts indicated some of the medications may be present for noninvasive therapies, but not as a continuous infusion (Supplementary Material). However, for the previously developed definition, only medication presence was used due to complexities in algorithmically processing medication information in relation to respiratory therapies. In future iterations, we would update the phenotype definition to include administration method rather than simply asking the model to look for concept presence. Finally, we propose development of industry standards for evaluation of LLM-based methods specifically with respect to automated processes to ensure appropriate oversight.
1.6. Conclusion
@@
[sorting=none]